# Rational torsion on hyperelliptic jacobian varieties

### v1

@@ 3. Construction of curves
Consider the following family of hyperelliptic curves of genus $3$
$C_{t}:\,y^{2}=(2(x^{2}+t)^{2}+(x^{2}+t)+1)^{2}+4(2(x^{2}+t)+1)$
over ${\mathbb{Q}}(t)$. Since the discriminant of $C_{t}$ is given by $2^{48}\cdot 5^{4}\cdot(1+t)(5+5t+4t^{3})$, if $t\neq-1$, then the class of the divisor $-(1:-2:0)$ ( 1 : - 2 : 0 ) on the Jacobian of $C_{t}$ is of order $10$.
@@ 4. Explicit families
There exists infinitely many hyperelliptic curves of genus $3$ over ${\mathbb{Q}}$ whose Jacobians posses a rational point of order $13$.
Proof: In Theorem 3.2, pick $r(x)=ax+b$, $a\neq 0$, to be a linear polynomial in ${\mathbb{Q}}[x]$, $u(x)=u\in{\mathbb{Q}}\setminus\{0\}$, and $a_{1}(x)=cx+d\in{\mathbb{Q}}[x]$, $c\neq 0$. When the discriminant of this curve is nonzero, we get a $5$-parametric family of hyperelliptic curves satisfying the hypothesis of the theorem, where that divisor at infinity is $(1:a^{2}\cdot c^{2}\cdot u:0)-(1:-a^{2}\cdot c^{2}\cdot u:0)$. $\Box$
@@
There exists infinitely many hyperelliptic curves of genus $4$ over ${\mathbb{Q}}$ whose Jacobians posses a rational point of order $15$.
@@
If $u\neq 0$, then the hyperelliptic curve $C^{13}_{u,t}$ of genus $3$ described by the equation
$y^{2}=(x+t)^{2}(2t+u+x+u(x+t)(x+2t)^{2})^{2}+4(t+x+u(x+t)^{2}(x+2t))$
satisfies that the divisor at infinity is torsion of order $13$.
Example 4.4.
Consider the genus $4$ hyperelliptic curve described by
$C^{15}_{s,t}:\,y^{2}=(x+t)^{2}\left(x+(x-s)(1+x^{2}(x+t))\right)^{2}+4\left(t+x+x(-s+x)(t+x)^{2}\right).$
For a choice of $s,t$ such that the discriminant of $C_{s,t}$ is nonzero, one sees that the divisor at infinity $-(1:-1:0)$ ( 1 : - 1 : 0 ) is torsion of order $15$.
Remark 4.5.
In examples 4.3, the equation describing the curve $C^{13}_{u,t}$ is of the form $(x+t)f_{u,t}(x)$ where the polynomial $f_{u,t}$ has generic Galois group $\textbf{S}_{7}$. The latter statement can be justified by seeing that the specialization $f_{1,1}$ has Galois group $\textbf{S}_{7}$. It follows that $f_{u,t}$ has Galois group $\textbf{S}_{7}$ for infinitely many pairs $u,t$. According to Theorem 3.2, this means that the Jacobian $J_{C^{13}_{u,t}}$ of $C^{13}_{u,t}$ satisfies that $\operatorname{End}(J_{C^{13}_{u,t}})={\mathbb{Z}}$ and that $J_{C^{13}_{u,t}}$ is an absolutely simple abelian variety for infinitely many values of $u,t$.
The same argument yields that the Jacobian of $C^{15}_{s,t}$ in Example 4.4 is an absolutely simple variety for infinitely many choices of $s,t$.
We notice that for $g\geq 5$, there are at least two integers $N$ in the interval $[3g,4g+1]$ when $g$ is odd, $[3g,4g-1]$ when $g$ is even, satisfying the hypotheses of Corollary 3.4.
Example 4.6.
Consider the following families of hyperelliptic curves of genus $5$
$C^{17}_{s,t}$ ::: $y^{2}=4(t+x+x(t+x)^{2}(x^{2}+s))+(x+t)^{2}(x+(x^{2}+s)(1+x^{2}(x+t)))^{2}$
$C^{18}_{s,t,u}$ ::: $y^{2}=(x^{2}+s)^{2}(t+u+x+u(x+t)^{2}(x^{2}+s))^{2}+4(s+x^{2}+u(x+t)(x^{2}+s)^{2})$
$C^{21}_{s,t,u}$ ::: $y^{2}=4(s+x+u(x+s)^{2}(x^{2}+t))+(x+s)^{2}(t+u+x^{2}+u(x+s)(x^{2}+t)^{2})^{2}$

### v2

@@ 3. Construction of curves
The following equation
$C_{t}:\,y^{2}=(2(x^{2}+t)^{2}+(x^{2}+t)+1)^{2}+8(x^{2}+t)+4$
describes a family of hyperelliptic curves of genus $3$ over ${\mathbb{Q}}(t)$. The discriminant of the latter equation is given by $2^{48}\cdot 5^{4}\cdot(1+t)(5+5t+4t^{3})$. If $t\in{\mathbb{Q}}\setminus\{-1\}$, then the class of the divisor (1:2:0)(1:2:0)(1:2:0)-(1:-2:0) on the Jacobian of $C_{t}$ is of order $10$.
@@ 4. Explicit families
There exists infinitely many non-isomorphic hyperelliptic curves of genus $3$ over $\mathbb{Q}$} whose Jacobians posses a rational point of order $13$.
Proof: In Theorem 3.2, pick $r(x)=ax+b$, $a\neq 0$, to be a linear polynomial in ${\mathbb{Q}}[x]$, $u(x)=u\in{\mathbb{Q}}\setminus\{0\}$, and $a_{1}(x)=cx+d\in{\mathbb{Q}}[x]$, $c\neq 0$. When the discriminant of this curve is nonzero, we get a $5$-parameter family of hyperelliptic curves of genus $3$ satisfying the hypothesis of the theorem, where that divisor at infinity is $(1:a^{2}\cdot c^{2}\cdot u:0)-(1:-a^{2}\cdot c^{2}\cdot u:0)$.
In order to show that there are infinitely many such non-isomorphic curves; we set $a=1$, $b=t$, $c=d=u=1$. This provides a $1$-parameter subfamily $C^{13}_{t}$ of the aforementioned $5$-parameter family of curves. In addition, the discriminant of the defining equation is $2^{28}h(t)$, where
$h(t)=36383+61280t-5762t^{2}-15163t^{3}+1662t^{4}+2470t^{5}-936t^{6}+49t^{7}+16t^{8}.$
In particular, if $t\in{\mathbb{Q}}$, then $C^{13}_{t}$ is a hyperelliptic curve. If two such curves $C^{13}_{t_{1}}$ and $C^{13}_{t_{2}}$, $t_{1},t_{2}\in{\mathbb{Q}}$, are isomorphic over $\mathbb{Q}$}, then $\Delta_{t_{1}}/\Delta_{t_{2}}$ is a $56$-th power of a non-zero rational number, [20, §2], where $\Delta_{t_{i}}$ is the discriminant of the defining equation of $C^{13}_{t_{i}}$, $i=1,2$. A classical application of the Sieve of Eratosthenes can be used to prove that the integers $n$ for which $h(n)$ is $8$-th power free have positive density, [7]. This implies that there are infinitely many non-isomorphic hyperelliptic curves in the family $C^{13}_{t}$. $\Box$
@@
There exists infinitely many non-isomorphic hyperelliptic curves of genus $4$ over $\mathbb{Q}$} whose Jacobians posses a rational point of order $15$.
Proof: In Theorem 3.2, we set $r(x)=a_{1}(x)=x+1$, $u(x)=x+t$. This gives rise to a $1$-parameter family $C_{t}^{15}$ of hyperelliptic curves of genus $4$ with a torsion divisor at infinity of order $15$. In addition, the discriminant of the defining equation is $2^{36}h(t)$ where
$h(t)$ $=$ $25815184+4208616t-12580992t^{2}+15705528t^{3}-6568623t^{4}-4422996t^{5}+6829050t^{6}-3728520t^{7}$
$+$ $1148643t^{8}-247912t^{9}+48114t^{10}-8748t^{11}+729t^{12}.$
Now we conclude as in the proof of Theorem 4.1. $\Box$
Remark 4.3.
In the proof of Theorem 4.1, the equation describing the curve $C^{13}_{t}$ is of the form $(x+t)f_{t}(x)$ where the polynomial $f_{t}$ has generic Galois group $\textbf{S}_{7}$. The latter statement can be justified by seeing that the specialization $f_{1}$ has Galois group $\textbf{S}_{7}$. Hilbert’s Irreducibility Theorem implies that the Galois groups of $f_{1}$ and $f_{t}$ over $\mathbb{Q}$} coincide for infinitely many rational values of $t$. According to Theorem 3.2, this means that the Jacobian $J_{C^{13}_{t}}$ of $C^{13}_{t}$ satisfies that $\operatorname{End}(J_{C^{13}_{t}})={\mathbb{Z}}$ and that $J_{C^{13}_{t}}$ is an absolutely simple abelian variety for infinitely many rational values of $t$.
The same argument yields that the Jacobian of $C^{15}_{t}$ in the proof of Theorem 4.2 is an absolutely simple variety for infinitely many rational values of $t$.
We notice that for $g\geq 5$, there are at least two integers $N$ in the interval $[3g,4g+1]$ when $g$ is odd, [3g,4g1][3g,4g-1] when $g$ is even, satisfying the hypotheses of Corollary 3.4.
@@
Consider the following families of hyperelliptic curves of genus 55
$C^{17}_{s,t}$ $:$ $y^{2}=4(t+x+x(t+x)^{2}(x^{2}+s))+(x+t)^{2}(x+(x^{2}+s)(1+x^{2}(x+t)))^{2}$
$C^{18}_{s,t,u}$ $:$ $y^{2}=(x^{2}+s)^{2}(t+u+x+u(x+t)^{2}(x^{2}+s))^{2}+4(s+x^{2}+u(x+t)(x^{2}+s)^{2})$
$C^{21}_{s,t,u}$ $:$ $y^{2}=4(s+x+u(x+s)^{2}(x^{2}+t))+(x+s)^{2}(t+u+x^{2}+u(x+s)(x^{2}+t)^{2})^{2}$